OpenStudy (anonymous):

witch point is represented on the line y + 4 = -5 (x-3)? need help

OpenStudy (knick032):

Point-Slope form:\[(y-y _{1}) = m(x-x _{1})\] So \[y _{1} = -4\] and \[x _{1}=3\]